# Few-Shot API Attack Anomaly Detection in a Classification-by-Retrieval Framework

### v1

@@ VII Experimental Results
Several observations stand out when considering the framework performance using BERT and RoBERTa embedding compared to FastText. It’s noticeable that BERT and RoBERTa introduce longer training times for all models. For instance, NMSLIB-COS with BERT takes around 0.446 seconds, while with RoBERTa, it’s approximately 0.563 seconds, compared to the original 0.0552 seconds with FastText. This increase in training time could be attributed to the more computational complexity nature of BERT [70, 71] and RoBERTa [72] models. In terms of performance, when considering the BERT embedding, NMSLIB-COS framework continues to exhibit robust performance, achieving an F1-score of 0.9675. This result suggests that the framework effectively leverages the contextual information embedded within BERT’s representations to identify anomalies. The high precision (0.947) and recall (0.9982) values further support the framework’s ability to maintain a fine balance between detecting true anomalies and minimizing false positives. However, a noteworthy observation lies in the performance of NMSLIB-COS framework when using RoBERTa embeddings. Surprisingly, while RoBERTa is considered an even more advanced and powerful language model compared to BERT, the F1-score for NMSLIB-COS drops slightly to 0.9664. This outcome raises questions about why the transition to RoBERTa, which typically exhibits superior performance across range of natural language processing tasks [67, 73, 74], did not lead to an improved performance for this specific outlier detection method. Generally, both BERT and RoBERTa maintain performance comparable with FastText with some exceptions. NMSLIB-COS with BERT achieves an F1-score of 0.9675, a slight improvement over FastText’s 0.9713. Similarly, GMM, MCD, CBLOF, and IFOREST also exhibit consistent or improved F1-scores with BERT and RoBERTa. However, for DEEPSVDD, the F1-scores drop slightly with BERT to 0.9187 and even more with RoBERTa to 0.8999. Precision and recall also showcase similar trends.
The ANN baseline models generally perform better than the traditional outlier detection methods, and their performance is quite similar with only slight differences. Among these models, HNSWLIB-IP stands out with the highest F1 score (0.9765) and perfect recall (100%), which means it doesn’t miss any actual anomalies. When we balance training time, precision, and recall, the BRUTEFORCE-BLAS models achieve the best precision and recall while needing less training time. This makes them suitable for scenarios where quick responses or limited resources are important. On the other hand, models like HNSWLIB-IP and ANNOY-MANHATTAN have slightly longer training times but achieve slightly higher F1 scores. The relatively slower performance of the MINKOWSKI distance metric in the BRUTEFORCE-BLAS model might be due to the more complicated calculations involved in the MINKOWSKI distance, especially when dealing with high-dimensional data [75]. The ANNOY model takes a different approach to constructing its search structure, which might explain its slightly slower training times compared to other models. ANNOY’s method involves generating random projections and building binary trees for its search index, which requires significant computation to ensure effective projection dimensions that preserve data relationships [76]. Despite not having the highest F1 score, NMSLIB finds use in industries like Amazon Elasticsearch Service [77]. This indicates that although its accuracy might not be the absolute best compared to other ANN implementations, its simple deployment and easy integration align well with the requirements of real-world applications. The overall performance of the algorithm can be evaluated through the average outcomes of difference distance metrics and the algorithm parameters. As depicted in Figure  7, each algorithm utilizing FastText embeddings exhibited significantly swifter processing times compared to BERT and RoBERTa. FastText and CKDTREE attained the shortest average index build times (0.00584 seconds), whereas BALLTREE operated over 2.5 times slower despite employing the same embeddings. Within the top 8 fastest FastText algorithms, only BRUTEFORCE demonstrated favorable performance when applied to BERT and RoBERTa, achieving scores of 0.02745 and 0.03215, respectively. Conversely, ANNOY’s performance was subpar for each language model. Regarding model accuracy, as illustrated in Figure  7, every algorithm displayed enhanced results when utilizing FastText, achieving F1-scores exceeding 97%. Similarly, the remaining algorithms also exhibited strong performance with F1-scores surpassing 96%. Generally, our framework achieved better accuracy compared to previous unsupervised and semi-supervised studies using the CSIC 2010 dataset, as shown in Table  I.
The results for the ATRDF 2023 dataset, as can be seen from Table III, show that we observed a perfect classifier for most of ANN algorithms. With 100% F1-score, recall and precision, our framework exceeded the traditional outlier detection baseline requirements. Most of the models demonstrate the shortest training/building times, whereas AUTOENCODER, MCD and DEEPSVDD show significantly longer train times of average 5.3368, 109.6675 and 3.3173 seconds respectively. While NMSLIB-COS did not attain the briefest build time, it can still be regarded as relatively rapid. It is evident that BERT and RoBERTa result in extended training durations for the majority of models, with a few exceptions. LMDD, CBLOF, QMCD, HBOS, and AUTOENCODER exhibited comparatively swifter training completion when utilized with BERT and RoBERTa. Most of the traditional outlier detection baselines also performed well but struggled to predict all positive classes, which affected the precision score. LMDD obtained an average 49.41% in F1-Score, which is much worse than those tested. QMCD failed to predict when tested with BERT and RoBERTa. The comprehensive mean benchmark outcomes of the diverse distance metrics and algorithm parameters are presented in Figure  8. FastText and CKDTREE managed to achieve the briefest average index construction times (0.00007 seconds), whereas BALLTREE functioned more than twice as slowly even though it utilized the same embeddings. Generally, all the algorithms exhibited rapid performance, with the exception of ANNOY across all language models. As previously mentioned, all ANN algorithms demonstrated outstanding predictive accuracy.
Train Time (Sec) F1-Score Precision Recall
BERT Fast Text Ro BERTa BERT Fast Text Ro BERTa BERT Fast Text Ro BERTa BERT Fast Text Ro BERTa
HNSWLIB-IP 4.4549 0.0978 3.6719 0.9443 0.9766 0.7548 0.8890 0.9547 0.6062 1.0000 1.0000 1.0000
ANNOY-MANHATTAN 11.426 3.4422 11.924 0.9690 0.9728 0.9683 0.9507 0.9576 0.9496 0.9974 0.9941 0.9973
FAISS 0.3592 0.0473 0.9377 0.9682 0.9725 0.9680 0.9493 0.9571 0.9491 0.9974 0.9940 0.9973
BRUTEFORCE-BLAS-COSINE 0.0191 0.0005 0.0179 0.9682 0.9725 0.9680 0.9493 0.9571 0.9491 0.9974 0.9941 0.9972
BRUTEFORCE-BLAS-CITYBLOCK 0.0132 0.0004 0.0122 0.9682 0.9724 0.9680 0.9493 0.9570 0.9491 0.9974 0.9941 0.9973
BRUTEFORCE-BLAS-MANHATTAN 0.0176 0.0004 0.0165 0.9682 0.9724 0.9680 0.9493 0.9570 0.9491 0.9974 0.9941 0.9973
KDTREE-CITYBLOCK 1.5269 0.0236 1.4969 0.9682 0.9724 0.9680 0.9493 0.9570 0.9491 0.9974 0.9941 0.9973
KDTREE-MANHATTAN 1.5325 0.0285 1.4807 0.9682 0.9724 0.9680 0.9493 0.9570 0.9491 0.9974 0.9941 0.9973
BALLTREE-CITYBLOCK 1.2906 0.0145 1.3317 0.9682 0.9724 0.9680 0.9493 0.9570 0.9491 0.9974 0.9941 0.9973
BALLTREE-MANHATTAN 1.2704 0.0141 1.3043 0.9682 0.9724 0.9680 0.9493 0.9570 0.9491 0.9974 0.9941 0.9973
CKDTREE 0.3944 0.0058 0.4961 0.9682 0.9723 0.9681 0.9493 0.9567 0.9492 0.9974 0.9942 0.9973
KDTREE-EUCLIDEAN 1.5244 0.0214 1.5159 0.9682 0.9723 0.9680 0.9493 0.9567 0.9491 0.9974 0.9942 0.9973
KDTREE-L2 1.5497 0.0250 1.5180 0.9682 0.9723 0.9680 0.9493 0.9567 0.9491 0.9974 0.9942 0.9973
KDTREE-MINKOWSKI 1.5287 0.0347 1.5336 0.9682 0.9723 0.9680 0.9493 0.9567 0.9491 0.9974 0.9942 0.9973
BALLTREE-EUCLIDEAN 1.2903 0.0154 1.2902 0.9682 0.9723 0.9680 0.9493 0.9567 0.9491 0.9974 0.9942 0.9973
BALLTREE-L2 1.2916 0.0153 1.2997 0.9682 0.9723 0.9680 0.9493 0.9567 0.9491 0.9974 0.9942 0.9973
BALLTREE-MINKOWSKI 1.2732 0.0154 1.3031 0.9682 0.9723 0.9680 0.9493 0.9567 0.9491 0.9974 0.9942 0.9973
BRUTEFORCE-BLAS-EUCLIDEAN 0.0108 0.0005 0.0116 0.9682 0.9722 0.9680 0.9493 0.9566 0.9491 0.9974 0.9942 0.9973
BRUTEFORCE-BLAS-MINKOWSKI 0.0765 0.1127 0.1025 0.9682 0.9722 0.9680 0.9493 0.9566 0.9491 0.9974 0.9942 0.9973
NMSLIB-L2 0.9779 0.0473 1.0642 0.9680 0.9721 0.9674 0.9491 0.9568 0.9482 0.9973 0.9936 0.9972
NMSLIB-L1 0.9580 0.0489 1.1008 0.9680 0.9721 0.9676 0.9491 0.9567 0.9488 0.9973 0.9937 0.9969
ANNOY-ANGULAR 13.711 4.1185 13.759 0.9672 0.9720 0.9684 0.9477 0.9564 0.9497 0.9973 0.9941 0.9973
ANNOY-EUCLIDEAN 12.177 4.5989 13.148 0.9680 0.9720 0.9686 0.9491 0.9563 0.9501 0.9974 0.9942 0.9973
HNSWLIB-L2 6.5040 0.6606 6.8610 0.9673 0.9719 0.9659 0.9479 0.9562 0.9458 0.9973 0.9939 0.9970
NMSLIB-COS (FT-ANN) 0.4458 0.0552 0.5626 0.9675 0.9713 0.9664 0.9470 0.9538 0.9465 0.9982 0.9954 0.9972
GMM 3.3663 0.1052 3.3310 0.9380 0.9651 0.9686 0.9290 0.9395 0.9476 0.9494 0.9931 0.9948
HNSWLIB-COSINE 6.4246 0.4612 7.2954 0.9462 0.9642 0.9433 0.9079 0.9414 0.9023 0.9994 0.9978 1.0000
MCD 1269.3 6.3603 1261.3 0.9622 0.9572 0.9517 0.9377 0.9382 0.9159 0.9971 0.9774 0.9985
CBLOF 2.8297 0.5747 2.7095 0.9404 0.9445 0.9365 0.9297 0.9318 0.9276 0.9531 0.9589 0.9477
IFOREST 16.066 0.5734 15.980 0.9397 0.9419 0.9362 0.9289 0.9323 0.9288 0.9525 0.9538 0.9464
HBOS 0.9967 0.8365 1.0126 0.9383 0.9393 0.9367 0.9289 0.9311 0.9274 0.9501 0.9502 0.9483
PCA 27.829 0.0575 28.413 0.9380 0.9391 0.9349 0.9290 0.9304 0.9271 0.9494 0.9504 0.9454
AUTOENCODER 82.873 75.064 84.402 0.9380 0.9389 0.9349 0.9290 0.9303 0.9271 0.9494 0.9502 0.9454
KDE 86.801 3.6710 83.525 0.9370 0.9361 0.9338 0.9283 0.9296 0.9263 0.9483 0.9460 0.9441
OCSVM 45.391 1.6392 50.025 0.9368 0.9360 0.9326 0.9283 0.9295 0.9255 0.9480 0.9459 0.9426
FB 10.875 10.547 10.775 0.9355 0.9354 0.9319 0.9279 0.9303 0.9255 0.9460 0.9444 0.9415
LOF 1.1774 1.0486 1.1147 0.9354 0.9350 0.9317 0.9279 0.9301 0.9255 0.9459 0.9439 0.9412
DEEPSVDD 38.214 32.719 38.242 0.9187 0.9235 0.8999 0.9230 0.9381 0.9155 0.9210 0.9126 0.8954
QMCD 9.9598 0.7783 10.251 0.0000 0.3500 0.0000 0.0000 0.8737 0.0000 0.0000 0.2259 0.0000
LMDD 11883 225.87 11230 0.3292 0.3017 0.3101 0.5701 0.5280 0.5885 0.2476 0.2293 0.2250
TABLE II: Performance comparisons of our framework (FT-ANN) on the CSIC 2010 dataset
Figure 7: Model level ANN benchmark on the CSIC 2010 dataset
Train time (Sec.) F1-score Precision Recall
BERT FAST TEXT Ro BERTa BERT FAST TEXT Ro BERTa BERT FAST TEXT Ro BERTa BERT FAST TEXT Ro BERTa
ANNOY-ANGULAR 0.9859 0.0908 1.0065 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
KDTREE-CITYBLOCK 0.0007 0.0001 0.0007 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
NMSLIB-L2 0.0026 0.0020 0.0025 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
KDTREE-EUCLIDEAN 0.0008 0.0002 0.0008 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
ANNOY-EUCLIDEAN 0.9808 0.0917 1.0144 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
KDTREE-L2 0.0008 0.0002 0.0008 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
KDTREE-MANHATTAN 0.0007 0.0001 0.0007 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
FAISS 0.0015 0.0010 0.0004 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
NMSLIB-COS (FT-ANN) 0.0039 0.0029 0.0042 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
CKDTREE 0.0005 0.0001 0.0005 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
KDTREE-MINKOWSKI 0.0007 0.0002 0.0008 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
BRUTEFORCE-BLAS-MINKOWSKI 0.0087 0.0028 0.0003 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
BRUTEFORCE-BLAS-MANHATTAN 0.0002 0.0001 0.0002 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
BRUTEFORCE-BLAS-EUCLIDEAN 0.0003 0.0001 0.0002 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
BRUTEFORCE-BLAS-COSINE 0.0002 0.0001 0.0002 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
BRUTEFORCE-BLAS-CITYBLOCK 0.0002 0.0001 0.0002 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
BALLTREE-MINKOWSKI 0.0005 0.0002 0.0005 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
BALLTREE-MANHATTAN 0.0005 0.0001 0.0005 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
BALLTREE-L2 0.0005 0.0001 0.0005 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
BALLTREE-EUCLIDEAN 0.0005 0.0002 0.0005 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
BALLTREE-CITYBLOCK 0.0005 0.0001 0.0005 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
NMSLIB-L1 0.0023 0.0019 0.0026 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
ANNOY-MANHATTAN 0.9944 0.0937 1.0303 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
HNSWLIB-L2 0.0010 0.0005 0.0010 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
DEEPSVDD 3.4848 3.1860 3.2811 0.9953 0.9960 0.9944 0.9908 0.9921 0.9889 1.0000 1.0000 1.0000
GMM 2.1158 0.1115 2.1114 0.9949 0.9953 0.9947 0.9899 0.9907 0.9896 1.0000 1.0000 1.0000
KDE 0.0043 0.0010 0.0044 0.9947 0.9946 0.9947 0.9896 0.9893 0.9896 1.0000 1.0000 1.0000
IFOREST 0.2578 0.1915 0.2597 0.9944 0.9946 0.9944 0.9888 0.9893 0.9888 1.0000 1.0000 1.0000
OCSVM 0.0028 0.0011 0.0028 0.9939 0.9946 0.9941 0.9879 0.9893 0.9884 1.0000 1.0000 1.0000
FB 0.0207 0.0144 0.0210 0.9944 0.9946 0.9944 0.9888 0.9893 0.9888 1.0000 1.0000 1.0000
PCA 0.0132 0.0015 0.0140 0.9951 0.9946 0.9951 0.9903 0.9893 0.9903 1.0000 1.0000 1.0000
AUTOENCODER 5.0431 6.0279 4.9394 0.9951 0.9946 0.9951 0.9903 0.9893 0.9903 1.0000 1.0000 1.0000
LOF 0.0016 0.0014 0.0016 0.9944 0.9946 0.9944 0.9888 0.9893 0.9888 1.0000 1.0000 1.0000
HBOS 0.4646 0.9127 0.4715 0.9938 0.9944 0.9947 0.9878 0.9888 0.9896 1.0000 1.0000 1.0000
MCD 151.30 3.6180 174.05 0.9939 0.9942 0.9886 0.9879 0.9885 0.9887 1.0000 1.0000 0.9898
CBLOF 0.0672 0.4408 0.0609 0.9949 0.9896 0.9957 0.9900 0.9795 0.9916 1.0000 1.0000 1.0000
HNSWLIB-COSINE 0.0011 0.0005 0.0011 0.9835 0.9785 0.9870 0.6119 0.8101 0.4654 1.0000 1.0000 1.0000
HNSWLIB-IP 0.0008 0.0004 0.0008 0.9627 0.9578 0.9866 0.4124 0.8958 0.5951 1.0000 1.0000 1.0000
QMCD 0.0379 0.5941 0.0431 0.0000 0.8946 0.0000 0.0000 0.9832 0.0000 0.0000 0.8240 0.0000
LMDD 0.3028 0.3387 0.3208 0.5641 0.2920 0.6261 1.0000 0.8237 1.0000 0.3982 0.1825 0.4628
TABLE III: Performance comparisons of our framework (FT-ANN) on the ATRDF 2023 dataset
Figure 8: Model level ANN benchmark on the ATRDF 2023 dataset
VIII Discussion and Conclusions
This paper suggests an innovative unsupervised few-shot anomaly detection framework that leverages a dedicated generic language model for API based on FastText embedding and uses ANN search in a Classification-by-retrieval approach. We showed that API attacks could be easily identified with no previous learning. To the best of our knowledge, this is the first work to utilize a Classification-by-retrieval framework based on the generalized approach of FastText embeddings combined with the approximate search to find anomalies in API traffic. We present a unique pre-processing technique to enhance input generalization and simplify API structure. This approach encompasses dividing input data into individual tokens, then constraining the vocabulary to a limited set of tokens. Consequently, the API structure becomes streamlined as the number of unique tokens diminishes, enabling input generalization and enables high detection accuracy even with minimal examples per class. We presented several state-of-the-art models for this task, performed a comparative analysis and demonstrated the best accuracy on the CSIC 2010 and ATRDF 2023 datasets. We showcased multiple cutting-edge models for this objective through two other widely adopted language models BERT and RoBERTa. Our comprehensive analysis encompassed benchmarking various ANN search algorithms, where we illustrated our models’ exceptional accuracy on the CSIC 2010 and ATRDF 2023 datasets. While we noted that our proposed dense cosine distance approach utilizing NMSLIB did not yield the top F1 score among ANN algorithms, its straightforward implementation and seamless integration make it a suitable choice for practical applications, particularly those like Amazon Elasticsearch Service that prioritize real-world compatibility. One notable limitation of this and similar studies arises from the scarcity of up-to-date and representative network traffic datasets. Many research efforts rely solely on the CSIC 2010 dataset, raising concerns about the reliability and applicability of such data, consequently impacting the quality and generalizability of traffic models. While our evaluation focuses on measuring the impact of FT-ANN on the HTTP dataset, it is important to acknowledge that APIs exist in diverse forms, including REST, GraphQL, gRPC, and WebSockets, each catering to specific use cases and carrying common vulnerabilities. Our future work entails expanding the evaluation to encompass a wider range of API types and to assess the performance across various loads and volumes.

### v2

@@ 6 Experimental Results
Several observations stand out when comparing the framework performance using BERT and RoBERTa embedding to FastText. Notably, BERT and RoBERTa result in longer training and testing times for all models. For instance, FT-ANN with FastText is approximately 87%-90% faster in training than RoBERTa and BERT, and similarly, around 78% faster in testing. This increase in time could be attributed to the more computational complexity nature of BERT (Yu et al., 2022; Xin et al., 2021) and RoBERTa (Rücklé et al., 2020) models. In terms of performance, when considering both RoBERTa and BERT embedding, FT-ANN continues to exhibit robust performance, achieving an F1-score of 0.9664 and 0.9675, respectively. This result suggests that the framework effectively leverages the contextual information embedded within the two language model representations to identify anomalies. The high precision and recall values for FT-ANN across all language models further support the framework’s ability to maintain a fine balance between detecting true anomalies and minimizing false positives. MCD with RoBERTa embedding achieves a recall of 0.9985, indicating better performance in correctly identifying positive instances compared to FT-ANN with RoBERTa embedding, which achieved 0.9972. Another noteworthy observation lies in the performance of the FT-ANN framework when using RoBERTa embedding. Surprisingly, while RoBERTa is considered an even more advanced and powerful language model compared to BERT, the F1-score for FT-ANN drops slightly to 0.9664. This outcome raises questions about why the transition to RoBERTa, which typically exhibits superior performance across a range of natural language processing tasks (Liu et al., 2019; Tarunesh et al., 2021; Rajapaksha et al., 2021), did not lead to an improved performance for this specific outlier detection method. MCD, KDE, OCSVM, FB, LOF, and LMDD exhibit consistent or improved F1-scores with BERT. However, for DeepSVDD, the F1-scores drop slightly with BERT to 0.9187 and even more with RoBERTa to 0.8999.
FT-ANN demonstrates superior performance over traditional outlier detection methods, while other ANN algorithms show similar performance with only minor variations. One of the most important challenges in any cybersecurity system is to minimize FP and take action in real time (Kabir and Hartmann, 2018). Therefore, we conducted a trade-off analysis between QPS and index build time (training) with precision. In Figures 7 and 8, we show only the leading algorithms for clarity, while the complete file is available in the project GitHub repository (Aharon et al., 2024). The results show that FT-ANN (Nmslib-COS) offers a compelling balance of speed and accuracy, with a QPS of 4795.43 and a precision of 0.9498, FT-ANN excels in delivering fast query responses while maintaining competitive precision. Hnswlib-IP achieves the highest QPS (6,065.70) and slightly higher precision (0.9547), and Nmslib-L1 and Nmslib-L2 also show strong performance with high QPS and slightly higher precision. Algorithms like Bruteforce-Blas-Cosine and Faiss, despite their highest precision, significantly lag in QPS. When considering build time, FT-ANN demonstrates a moderate build time of 0.0519 seconds. While some algorithms like Bruteforce-Blas variants have significantly shorter build times (e.g., Bruteforce-Blas-Cityblock with 0.0004 seconds), they lag far behind in QPS. On the other hand, algorithms like Hnswlib-Cosine and Hnswlib-L2, despite achieving relatively high QPS, have longer build times (0.461 and 0.660 seconds, respectively). The Annoy algorithm’s slower build time and lower QPS may be due to generating random projections and building binary trees for its search index, which requires significant computation to maintain data relationships within projection dimensions, leading to slightly longer training times compared to other models (Cheng et al., 2021).
Figure 7: Precision-QPS (queries per second) trade-off for top-performing in-memory ANN algorithms on the CSIC 2010 dataset — higher and to the right indicates better performance.
Figure 8: Precision-Index build time trade-off in seconds for top-performing in-memory ANN algorithms on the CSIC 2010 dataset — lower and to the right indicates better performance.
The results with BERT and RoBERTa embedding show significant differences in the performance metrics. FT-ANN with BERT achieves a precision of 0.9472, a QPS of 2,189.39, and a build time of 1.0195 seconds, while with RoBERTa, it attains a precision of 0.9484, a QPS of 1,591.54, and a build time of 1.2153 seconds. Other algorithms like Hnswlib-IP and Nmslib-L1 show decreased QPS and longer build times with BERT and RoBERTa embedding, reflecting the increased computational complexity and processing requirements of these advanced models. Despite maintaining high precision, algorithms such as Faiss and Bruteforce-Blas variants exhibit significantly lower QPS, highlighting their limitations in real-time applications. These insights underline that while BERT and RoBERTa embedding provide high precision, they often come at the cost of reduced speed and increased computational demands compared to FastText.
Despite not having the highest accuracy, Nmslib is used in industries such as Amazon Elasticsearch Service (Amazon Web Services, 2020). This indicates that although its accuracy might not be the absolute best compared to other ANN implementations, its simple deployment, ease of integration, and excellent balance between speed and accuracy align well with the requirements of real-world applications.
As seen in Table 4, FT-ANN demonstrates superior accuracy compared with current SOTA unsupervised learning models on the CSIC 2010 dataset, specifically AE variants (Mac et al., 2018; Demirel and Sandikkaya, 2023; Vartouni et al., 2018; Moradi Vartouni et al., 2019) and SVDD (Vartouni et al., 2023).
The results for the ATRDF 2023 dataset, as can be seen from Table 5, show that FT-ANN consistently achieves perfect scores across all metrics (F1 score, precision, recall) with BERT, FastText, and RoBERTa embedding. Comparatively, most of the baseline models also achieved high accuracy but failed to classify correctly all positive instances as reflected in the precision. DeepSVDD, AE, and IF perform well but exhibit longer build and test times. LMDD has higher build and test times, indicating again less efficient scalability.
Most ANN algorithms perform excellently on ATRDF 2023, so we chose to focus on comparing QPS and build time across different methods to evaluate their speed and scalability. The results are shown in Figures 9 and 10. FT-ANN using FastText embedding demonstrates a strong balance with high QPS (10,729.09) and reasonable build time (0.002997 seconds). In comparison, Nmslib-L1 and Nmslib-L2 outperform FT-ANN in terms of QPS, with values of 11,863.19 and 11,500.90, respectively, and slightly lower build times. Faiss and Hnswlib variants also offer high QPS with minimal build time. In contrast, BallTree variants achieve modest QPS with extremely low build times. Bruteforce-Blas models, despite their simplicity and extremely low build times, achieve lower QPS, making them less suitable for high-throughput applications but beneficial where build time is critical. Generally, we observe the consistency in the model’s performance relative to each other under the different language model limitations.
Method F1 Score Precision Recall
FT-ANN 0.9713 0.9538 0.9954
Mac et al. (2018) 0.9463 0.9464 0.9462
Demirel and Sandikkaya (2023) 0.9274 0.9352 0.9333
Vartouni et al. (2023) 0.89 0.8583 0.9243
Vartouni et al. (2018) 0.8412 0.8029 0.8832
Moradi Vartouni et al. (2019) 0.8196 0.8122 0.8273
Table 4: Evaluation of CSIC 2010 dataset with various unsupervised methods.
F1 Score Precision Recall Train Time (Sec) Test Time (Sec)
BERT FastText RoBERTa BERT FastText RoBERTa BERT FastText RoBERTa BERT FastText RoBERTa BERT FastText RoBERTa
FT-ANN 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 0.0039 0.0029 0.0042 0.0012 0.0004 0.0012
DeepSVDD 0.9953 0.996 0.9944 0.9908 0.9921 0.9889 1.0 1.0 1.0 3.4848 3.186 3.2811 0.0383 0.0332 0.0476
AE 0.9951 0.9946 0.9951 0.9903 0.9893 0.9903 1.0 1.0 1.0 5.043 6.0278 4.9394 0.0421 0.0512 0.0412
FB 0.9944 0.9946 0.9944 0.9888 0.9893 0.9888 1.0 1.0 1.0 0.0207 0.0144 0.021 0.0139 0.0055 0.0146
IF 0.9944 0.9946 0.9944 0.9888 0.9893 0.9888 1.0 1.0 1.0 0.2578 0.1915 0.2597 0.0649 0.0307 0.0651
KDE 0.9947 0.9946 0.9947 0.9896 0.9893 0.9896 1.0 1.0 1.0 0.0043 0.001 0.0043 0.0068 0.0005 0.0069
LOF 0.9944 0.9946 0.9944 0.9888 0.9893 0.9888 1.0 1.0 1.0 0.0015 0.0014 0.0016 0.0015 0.0006 0.0011
OCSVM 0.9939 0.9946 0.9941 0.9879 0.9893 0.9884 1.0 1.0 1.0 0.0028 0.0011 0.0028 0.0017 0.0002 0.0016
PCA 0.9951 0.9946 0.9951 0.9903 0.9893 0.9903 1.0 1.0 1.0 0.0132 0.0014 0.014 0.0026 0.0002 0.0028
HBOS 0.9938 0.9944 0.9947 0.9878 0.9888 0.9896 1.0 1.0 1.0 0.4646 0.9126 0.4715 0.005 0.0228 0.0049
MCD 0.9939 0.9942 0.9886 0.9879 0.9885 0.9887 1.0 1.0 0.9898 151.32 3.618 174.05 0.2682 0.0005 0.3775
CBLOF 0.9949 0.9896 0.9957 0.99 0.9795 0.9916 1.0 1.0 1.0 0.0672 0.4408 0.0609 0.002 0.0006 0.002
LMDD 0.5641 0.292 0.6261 1.0 0.8237 1.0 0.3982 0.1825 0.4628 0.3028 0.3387 0.3208 1.2767 0.068 1.3278
Table 5: Performance evaluations of FT-ANN on the ATRDF 2023 dataset in comparison to baseline models for detecting outlying objects in multivariate data.
Figure 9: Evaluation of Index Build Time (seconds) for In-Memory ANN Algorithms on the ATRDF 2023 Dataset. The Y-axis represents algorithm names, and the X-axis shows build time in log scale.
Figure 10: Evaluation of QPS (queries per second) for In-Memory ANN Algorithms on the ATRDF 2023 Dataset. The Y-axis represents algorithm names, and the X-axis shows QPS in log scale.
7 Discussion and Conclusions
This paper suggests an innovative unsupervised few-shot anomaly detection framework that leverages a dedicated generic language model for API based on FastText embedding and uses ANN search in a Classification-by-retrieval approach. We demonstrated that API Injection attacks can be identified using only a few examples of normal API requests. Our framework not only accurately classifies the API traffic but also does it quickly and effectively with the support of incremental learning by the ANN, which is crucial for real-world systems. To the best of our knowledge, this is the first work to utilize a Classification-by-retrieval framework based on the generalized approach of FastText embedding combined with the approximate search to find anomalies in API traffic. Additionally, we presented a unique pre-processing technique to enhance input generalization and simplify API structure. This approach encompasses dividing input data into individual tokens and then constraining the vocabulary to a limited set of tokens. Consequently, the API structure becomes streamlined as the number of unique tokens diminishes, enabling input generalization and achieving high detection accuracy even with minimal examples per class. We presented several advanced anomaly detection models for this objective through two other widely adopted language models, BERT and RoBERTa. Our analysis also involved comparing various ANN search algorithms to demonstrate the framework’s performance on the CSIC 2010 and ATRDF 2023 datasets. We compared our framework to the current SOTA unsupervised techniques and showed that it achieved the highest accuracy on the CSIC 2010 dataset.
While we observed that the cosine similarity implementation in Nmslib doesn’t produce the highest F1 score among ANN algorithms, its straightforward implementation and seamless integration, and a good balance of QPS and build time make it a suitable choice for practical applications, particularly those like Amazon Elasticsearch Service that prioritize real-world compatibility.
One notable limitation of this and similar studies arises from the scarcity of up-to-date and representative network traffic datasets. Many research efforts rely solely on the CSIC 2010 dataset, raising concerns about the reliability and applicability of such data, consequently impacting the quality and generalizability of the resulting models. While our evaluation focuses on measuring the impact of FT-ANN on the HTTP dataset, it is important to acknowledge that APIs exist in diverse forms, including REST, GraphQL, gRPC, and WebSockets, each catering to specific use cases and carrying common vulnerabilities. Our future work entails expanding the evaluation to encompass a wider range of API attacks and types. Additionally, we intend to assess the stability of the framework and its resource utilization under varying loads and volumes.
Funding
This research received no specific grant from any funding agency in the public, commercial, or not-for-profit sectors.
CRediT authorship contribution statement
Udi Aharon: Conceptualization, Data curation, Formal analysis, Investigation, Methodology, Project administration, Resources, Software, Validation, Visualization, Writing – original draft, Writing– review & editing. Ran Dubin: Conceptualization, Methodology, Project administration, Supervision, Writing – review & editing. Amit Dvir: Conceptualization, Investigation, Methodology, Writing – review & editing. Chen Hajaj: Conceptualization, Data curation, Investigation, Methodology, Project administration, Writing – review & editing.